OpenStudy (anonymous):

I need to ask a question that contains an indefinite integral and theta, but only the "post" button is available to me. Where is the symbol button?

OpenStudy (unklerhaukus):

\[\int\theta \cdot \text d \theta\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\theta \cos \theta d \Theta u=\theta dv=\cos \Theta d \Theta du=d \Theta v=\sin \Theta\] I got: \[\int\limits_{}^{}\Theta \cos Thetad \Theta=\sin ^2\Theta+\cos \Theta+C\] Is this correct? Thanks:)